OpenStudy (jacksonjrb):

2x^(2/3)-x^(1/3)-1=0

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ 2x^{\LARGE\bf ⅔}-x^{\LARGE\bf ⅓}-1=0\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ 2\sqrt[3]{x^2}-\sqrt[3]{x}-1=0\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ 2\sqrt[3]{x^2}-\sqrt[3]{x}=1\LARGE\color{white}{ \rm │ }}\) I think then you would cbe troot both sides.

OpenStudy (solomonzelman):

*cube root both sides.

OpenStudy (mokeira):

would you cube root or cube to remove the roots?

OpenStudy (solomonzelman):

yeah, I mean to raise both sides to the 3rd power..

OpenStudy (jacksonjrb):

can you tell me what x would be equal to?

OpenStudy (solomonzelman):

yes, it is 1

OpenStudy (jacksonjrb):

if you plug 1 into the equation as x, it's not equal to 0

OpenStudy (solomonzelman):

yes it is

OpenStudy (solomonzelman):

2x^(2/3)-x^(1/3)-1=0 2(1)^(2/3)-(1)^(1/3)-1=0 2(1)-(1)-1=0 2 - 1 - 1 =0 0=0

OpenStudy (jacksonjrb):

ah okay can you explain how i get x equals 1 after i raise each side to the third power ?

OpenStudy (solomonzelman):

whatever you do, you don't change the value of the sides... right ?

OpenStudy (phi):

notice if you let y= x^(⅓) then \[ 2x^{⅔}-x^{⅓}-1=0 \\ 2y^2 -y -1 = 0 \] solve for y you can use the quadratic formula, or you can factor to find (2y +1)(y -1) =0 you get two answers for y. you can use them in y= x^(⅓) so x=y^3 to get the values for x

OpenStudy (jacksonjrb):

Thanks Phi. x is equal to -1 and +1

OpenStudy (phi):

no, rethink that

OpenStudy (jacksonjrb):

(2y +1)(y -1) =0 2y+1=0 y-1=0 y+1=0 y-1=0 y=-1 y=1 x=y^3 x=1^3 x=-1^3 x=1 x=-1

OpenStudy (jacksonjrb):

They both work if you plug them into the equation

OpenStudy (phi):

(2y +1)(y -1) =0 2y+1=0 or y-1=0 the idea is if you multiply two "things" and you get zero, then one or the other has to be zero y-1=0 add +1 to both sides and you get y= 1 now solve 2y + 1 = 0 the first step is add -1 to both sides. can you finish ?

OpenStudy (phi):

2y= -1 now divide both sides by 2

OpenStudy (jacksonjrb):

y=-1/2 x=-0.125

OpenStudy (phi):

so your two solutions are x=1 and x= -⅛ notice if we use x= -1 in the original equation then the cube root of x is -1 (because -1*-1*-1= -1) x^(⅓) = -1 and if we square it, i.e. x^(⅔) = -1*-1= 1 use those numbers in 2 x^(⅔) - x^(⅓) - 1 to get 2* 1 - (-1) - 1 2 + 1 - 1 2 we do not get zero... x=-1 is not a solution

OpenStudy (phi):

if x = -⅛ what is the cube root of x ?

OpenStudy (jacksonjrb):

Ah okay

OpenStudy (jacksonjrb):

0.5?

OpenStudy (phi):

½ * ½*½= ⅛ we want -⅛

OpenStudy (phi):

notice -½ * -½ * -½ works (= -⅛) so -½ is the cube root of -⅛ and if we square it to find x^(⅔) we get +¼

OpenStudy (jacksonjrb):

-1/2

OpenStudy (phi):

now use x^⅔ = ¼ and x^⅓ = -½ in the original

OpenStudy (jacksonjrb):

You get. 0

OpenStudy (phi):

yes

OpenStudy (jacksonjrb):

Mind helping me with one more similar problem? I already got it started

OpenStudy (phi):

please make it a new post.

OpenStudy (jacksonjrb):

Okay.